the ratio of the length and width of a rectangle is 3:8. if the length is 24 what is the width?

Answer by checkley79(3341)

3/8=24/N
3N=8*24
3N=192
N=192/3
N=64 ANS. FOR THE WIDTH
PROOF:
3/8=24/64 CROSS MULTIPLY
3*64=8*24
192=192